# Towards Building Multilingual Language Model for Medicine

### v2

@@ C.3 Detailed Results on Rating Scores
[b]
Method English Chinese Japanese French Russian Spanish Avg.
MMedLM 4.12 4.74 4.44 4.44 4.22 4.36 4.39
InternLM 3.58 4.04 3.26 2.74 2.64 2.8 3.18
BLOOMZ 2.68 3.18 3.02 3.2 3.28 3.22 3.1
LLaMA 2 3.24 3.2 3.46 3.2 4.18 3.72 3.5
PMC-LLaMA 4.3 3.22 3.06 3.22 3.26 3.2 3.38
GPT-3.5 3.08 2.62 3.76 4.2 3.42 3.7 3.46
[b] Method English Chinese Japanese French Russian Spanish Avg. MMedLM 4.3 4.78 4.94 4.14 4.44 4.66 4.54 InternLM 3.68 3.94 3.78 3.44 2.92 2.62 3.4 BLOOMZ 3.2 3.6 3.16 3.3 3.34 3.26 3.31 LLaMA 2 3.14 3.48 3.52 3.5 3.68 3.96 3.55 PMC-LLaMA 3.44 2.72 2.82 3.2 3.38 3.16 3.12 GPT-3.5 3.24 2.48 2.78 3.42 3.24 3.34 3.08
Appendix D Case Study
@@
D.1 Zero-shot Setting
In a zero-shot setting, we observe that it is challenging for a 7B-scale language model to consistently follow instructions, making it difficult to calculate accuracy automatically. Though difference cases, we notice that even when given the same instruction of "Answer with the best option directly.", the model will still express it in many ways. ###The most appropriate next step in management for this patient is to perform a bone marrow biopsy. (…) The best treatment for this patient is option C. (…) Given the patient’s symptoms and medical history, the most likely diagnosis is HIV infection. (…) The most appropriate next step in management for this patient is D. Ruxolitinib. (…)
What’s more, we find that in a zero-shot setting, the model is very sensitive to the instructions. For instance, only changing the instruction to "Answer the letter of the option directly" may lead the model to generate a different answer. Considering the difficulty in finding a fair way to automatically evaluate models’ performance in zero-shot setting, we turn to fine-tuning settings. The most appropriate next step in management for this patient is D. Ruxolitinib. This is because the patient’s symptoms (…) B### Explanation:The patient’s symptoms and laboratory findings suggest a myeloproliferative neoplasm (…)
D.2 Full Fine-tuning Setting
We further specifically compared the responses of MMedLM and InternLM, noting that MMedLM represents an advanced version of InternLM, augmented with domain-specific knowledge. As illustrated in Figures 6, 7, 8, 9, 10, and 11, MMedLM consistently demonstrates superior performance in selecting the correct option, unlike InternLM. Furthermore, MMedLM frequently justifies its choices with accurate reasons. For instance, in Figure 6, MMedLM accurately diagnoses ‘eosinophilic infiltration in the renal biopsy,’ subsequently applying its domain knowledge to identify these findings as indicative of toxic tubulointerstitial nephritis, leading to a precise diagnosis.
Nevertheless, we also notice an exception in Figure 10, where the language model (LLM) correctly identifies the answer, but with an erroneous rationale. Although the model attributes apoptosis to viral infections by stating that ’viruses attack cells, causing apoptosis,’ it overlooks the fact that apoptosis is, in reality, an active defense mechanism employed by the body to curb viral replication. This observation suggests that the model’s selection of the correct answer could be attributed to chance rather than a genuine understanding or possession of relevant knowledge. It further underscores the importance of evaluating not only the accuracy of the model’s answers but also its reasoning process.

### v3

@@ E.3 Detailed Results on Rating Scores
Appendix F Case Study
@@
F.1 Zero-shot Setting
In a zero-shot setting, we observe that it is challenging for a 7B-scale language model to consistently follow instructions, making it difficult to calculate accuracy automatically. Though difference cases, we notice that even when given the same instruction of "Answer with the best option directly.", the model will still express it in many ways. {mdframed}[backgroundcolor=gray!20]
  • ###The most appropriate next step in management for this patient is to perform a bone marrow biopsy. (…)
  • The best treatment for this patient is option C. (…)
  • Given the patient’s symptoms and medical history, the most likely diagnosis is HIV infection. (…)
  • The most appropriate next step in management for this patient is D. Ruxolitinib. (…)
What’s more, we find that in a zero-shot setting, the model is very sensitive to the instructions. For instance, only changing the instruction to "Answer the letter of the option directly" may lead the model to generate a different answer. Considering the difficulty in finding a fair way to automatically evaluate models’ performance in zero-shot setting, we turn to fine-tuning settings. {mdframed}[backgroundcolor=gray!20]
  • The most appropriate next step in management for this patient is D. Ruxolitinib. This is because the patient’s symptoms (…)
  • B### Explanation:The patient’s symptoms and laboratory findings suggest a myeloproliferative neoplasm (…)
F.2 Full Fine-tuning Setting
We specifically compared the responses of MMed-Llama 3 and Llama 3, noting that MMed-Llama 3 represents an advanced version of Llama 3, augmented with domain-specific knowledge. As illustrated in Supplementary Figures 1, 2, 3, 4, 5, and 6, MMed-Llama 3 demonstrates superior performance in selecting the correct answer, especially in multilingual cases. Moreover, MMed-Llama 3 can give more adequate reasons for its choice. For instance, in Supplementary Figure 1, MMed-Llama 3 accurately diagnoses the presence of ‘eosinophilic infiltration’ and ‘diffuse parenchymal inflammation on renal biopsy’, subsequently applying its domain knowledge to identify these findings as indicative of ‘tubulointerstitial nephritis’, leading to a precise diagnosis.